OpenStudy (anonymous):

Jacky has 3 cards for every 2 that Dorian owns. Together they have 280 cards. How many cards does Dorian have?

OpenStudy (ajprincess):

Let the number of cards Dorian has be x. Then number of cards Jack has=3x/2 3x/2+x=280 see if u can solve it nw.

OpenStudy (anonymous):

For every 3 cards of Jacky, Dorian has 2 cards.. So, Jacky has 3/5 of Total Cards = 3/5 of 280 = 168 And Dorian will have 2/5 of the total cards = 2/5 of 280 = 112.. So Dorian will have 112 cards...